A sample ofnicotine, weighs 6.899 grams. Convert this moles. Nicotine has the chemical formula C10H14N2. Be sure to include a unit with your answer.
Answer #1

Molecular formula of Nicotine is C10H14N2

So molecular weight of Nicotine= 10*12g + 14*1g + 2*14 = 120+14+28= 162 g

Weight of Nicotine sample = 6.899 g

Number of moles of Nicotine = 6.899 g / 162 g/mol = 0.0425 mol